A boy pull a wagon with a Force of 45N by means of a rope that makes an angle of 40 degree with the ground.how much work does he do in moving the wagon.
1
Expert's answer
2019-06-03T10:41:02-0400

Mechanical work

"W=F\cdot S\cdot cos(\alpha)"

We can define work per length

"\frac{W}{S}=F\cdot cos(\alpha)"

Calculating

"\frac{W}{S}=45\cdot cos(40)=34.5 (\frac{J}{m})"